OpenStudy (anonymous):

Solve 5x2 = –30x – 65

OpenStudy (anonymous):

\[x=-3\pm2i\] imaginary roots for this one

jhonyy9 (jhonyy9):

5x2 = -30x-65 5x2+30x+65=0 /5 x2 +6x +13=0 x1_2=(-6+/- sqrt(36-52))/2 = (-6+/- sqrt(-16))/2 = (-6+/- 4i)/2 = =2(-3+/- 2i)/2 = -3+/- 2i